40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 6 10:04:58 2014, manasa, Update, General, Unexpected power shutdown 
    Reply  Mon Oct 6 10:23:43 2014, Steve, Update, VAC, Unexpected power shutdown 
       Reply  Tue Oct 7 16:55:16 2014, Steve, Update, VAC, Unexpected sweaty valves sweatyV5.jpgsweatyV5f.jpg
          Reply  Wed Oct 8 08:46:57 2014, Steve, Update, VAC, VAT valves actuator lubricant 
             Reply  Mon Oct 13 14:44:52 2014, Steve, Update, VAC, RGA scan pd78 -day 13 RGAscan13Oct2014.png
                Reply  Mon Dec 1 17:13:57 2014, Steve, Update, VAC, RGA scan pd78 -day 63 RGAscanD63.png
                   Reply  Tue Mar 31 09:03:32 2015, Steve, Update, VAC, RGA scan pd78 day 182 RGApd78d182.png
                      Reply  Wed Jul 1 10:01:25 2015, Steve, Update, VAC, RGA scan pd78 day 275 rga275d.png
                         Reply  Mon Oct 5 10:11:35 2015, Steve, Update, VAC, RGA scan pd78 day 386 RGA_background.png
       Reply  Tue Dec 23 14:33:24 2014, Steve, Update, VAC, Chiara gets UPS prepUP.png
    Reply  Mon Oct 6 10:28:18 2014, manasa, Update, General, Unexpected power shutdown 
       Reply  Mon Oct 6 11:09:52 2014, Jenne, Update, General, Unexpected power shutdown: slow computers 
       Reply  Mon Oct 6 17:04:51 2014, ericq, Update, General, Unexpected power shutdown 
          Reply  Mon Oct 6 17:36:17 2014, ericq, Update, General, Unexpected power shutdown 
          Reply  Mon Oct 6 18:15:12 2014, manasa, Update, General, Unexpected power shutdown: end green alignment 
    Reply  Tue Oct 7 16:17:15 2014, Steve, Update, General, reason for 8 sec power outage 
Message ID: 10571     Entry time: Mon Oct 6 17:04:51 2014     In reply to: 10569     Reply to this: 10572   10573
Author: ericq 
Type: Update 
Category: General 
Subject: Unexpected power shutdown 

I brought back the PMC, MC and Arms.

PMC:

  • Same as when we replaced the busted sorensen, the kepco regulators in 1X1 (which power the FSS HV amp, PMC PZT and WFS) needed to be brought back up in the proper order. (Middle two are  + and - for the FSS, need to be rolled up in unison). Also the same as that occasion, sticky sliders prevented the full voltage range on the PMC PZT from being accessible. I touched every button on the PMC and FSS screens, which seemed to fix it. 
  • I then realigned the PMC to ~0.80 transmission

MC:

  • Needed to do some hand alignment to get a lock
  • Measured spot positions, they were all under 2mm
  • Despite centering the beams on the WFS and setting the offsets, WFS would not turn on successfully
  • Also, the autolocker on megatron isn't doing anything but blinking
  • Also also, MC2 is exhibiting some intermittent alignment wandering. The SUSDOF traces look like flat ramps lasting a few minutes. 

Arms:

  • No green was evident anywhere, but it didn't take to much alignment tweaking to get IR flashes
  • No signals were evident on RFPDs, confirmed light on PDs and power to demod boards. 
  • Turned out the 11MHz Marconi was not doing anything, and needed to be reset to the modulation frequency in ELOG 10314 (which reminds me that I need to update the sticker on the marconi)
  • Locked arms, ASS'ed, oplev spots were acceptable. 

 

ELOG V3.1.3-